# Adaptation of Agentic AI

### v2

@@ 5.2.3 Agentic Memory and Others
Memento [22] demonstrates that an agent’s memory system can be optimized as an external tool without any modification to the LLM planner. The system combines a frozen GPT-4.1 high-level planner with a trainable episodic case memory module. The memory stores past problem-solving trajectories, and the tool being trained is a neural Q-function that learns a case retrieval policy: which past cases to present to the frozen planner when facing a new problem. The training signal is remarkably simple: binary task success or failure. This sparse, trajectory-level reward is broadcast to all case-selection decisions in that trajectory, and a soft Q-learning algorithm updates the retrieval policy. Crucially, the frozen LLM never sees the Q-values or policy internals; it simply receives retrieved cases as context and generates its plan. Memento achieves top-tier performance: 87.88% on GAIA validation (ranked 1st), 79.40% on GAIA test (3rd place), and 95.0% on SimpleQA. Ablations show that case-based memory adds 4.7–9.6% absolute improvement on out-of-distribution tasks. This is remarkable because only the memory is trained; the same frozen LLM that performed worse without memory now excels simply because its information diet has been optimized.
Test-Time Memory Curation
@@
Another prominent example of T2 memory adaptation at inference-time is Dynamic Cheatsheet (DC) [185], a lightweight framework that provides a “persistent, evolving memory” for black-box LMs. The system operates “without modifying their underlying parameters” and requires no gradient-based updates. The framework consists of two core modules: a Solution Generator and a Memory Curator. This Memory Curator is the adaptive T2 tool. It operates without access to ground-truth labels ; instead, it has to assess the correctness and efficiency of the solutions by itself after they are produced by the frozen generator. Based on this self-assessment, the curator updates the memory by storing concise, transferable snippets such as “reusable strategies, code snippets, and general problem-solving insights” , rather than full, uncurated transcripts. ReasoningBank [186] extends this test-time curation concept by creating a memory framework that explicitly distills generalizable reasoning strategies from both successful and self-judged failed experiences. Unlike methods that store raw trajectories or only successful routines, ReasoningBank analyzes failures to extract “crucial preventative lessons”. This curated bank of reasoning strategies is then retrieved to guide the agent in future tasks. The framework also introduces memory-aware test-time scaling, which uses the curated memory to guide a scaled exploration, creating a “synergy between memory and test-time scaling” where the diverse experiences from scaling help forge stronger, more generalizable memories. This approach was shown to be effective on complex benchmarks like WebArena [187] and SWE-Bench [188].
Adapting the Embedding Space
@@
Training uses supervised learning on parallel sequences where ground-truth tool calls are replaced with toolken placeholders. The loss is masked so that only the toolken predictions (and subsequent argument tokens) contribute gradients to $W_{\tau}$. This is remarkably parameter-efficient: adding 234 tools requires training only $234\times 4096\approx 1M$ parameters (for LLaMA’s 4096-dimensional embeddings), compared to the 13B+ parameters of the full model. ToolkenGPT achieves 73% one-hop accuracy on FuncQA (vs. 57% for ReAct), 75% supervised accuracy on 234-relation KAMEL, and 68% success on VirtualHome with 58 action/object tools. Crucially, new tools can be added by simply expanding $W_{\tau}$ and continuing training—no full model retraining required. The conceptual contribution of ToolkenGPT is showing that adaptation can occur at the interface layer (the embedding space) rather than the parameter layer (the LLM weights), offering a middle ground between fully frozen T1 systems and fully fine-tuned A1/A2 systems.
Beyond the paradigms above, a diverse set of recent approaches further extends the tool adaptation framework. These methods introduce new training objectives, modalities, and architectural innovations that broaden the scope of tool adaptation.
@@
[…]
@@ 6 Comparison of Adaptation Paradigms
This section provides a comprehensive comparison of the four adaptation paradigms: (A1) Agent Adaptation with Tool Execution Signal, (A2) Agent Adaptation with Agent Output Signal, (T1) Agent-Agnostic Tool Adaptation, and (T2) Agent-Supervised Tool Adaptation. We first establish a conceptual framework for comparison, then analyze the agent-centric (A1/A2) and tool-centric (T1/T2) paradigms in depth, with special focus on the emergent “subagent-as-tool” and “graduation” concepts, and conclude with a quantitative synthesis of critical trade-offs.
@@ 6.1 A Framework for Comparison
  • Cost and Flexibility: We use cost to refer to compute and engineering effort required for adaptation, and flexibility to mean how easily the system’s behavior can be reconfigured. A1/A2 provide high parametric flexibility (the entire agent policy can change), whereas T1/T2 provide high system-level flexibility (capabilities can be added, swapped, or composed via tools) but remain bounded by the frozen agent’s intrinsic reasoning power.
  • Data Efficiency: Beyond raw compute, the amount of training data required differs dramatically across paradigms. Recent evidence suggests that T2 methods can match or surpass A2-style end-to-end agent training with orders of magnitude less data, by only training small subagents around a frozen backbone.
@@
  • Modularity and System Evolution: This axis focuses on engineering implications: how easily a system can be extended or maintained over time. Tool-centric paradigms (T1/T2) enable modular evolution and hot-swapping of components; agent-centric paradigms (A1/A2) tend to be monolithic and may suffer from catastrophic forgetting when adapted repeatedly.
In prose, the picture is as follows. Agent adaptation (A1/A2) is expensive but gives fine-grained control over the agent itself: one can rewrite the entire policy, alter reasoning style, alignment, and domain knowledge in a single model. This is high parametric flexibility, but each change typically requires retraining a large model and may unintentionally affect other behaviors. Tool adaptation (T1/T2), in contrast, does not touch the core agent; instead, it achieves system-level flexibility by letting us attach specialized tools, each tuned independently. We can freely augment separate capabilities (e.g., retrieval, planning, memory, code search), orchestrate them, and retire or replace tools without destabilizing the base agent. The trade-off is that these tools cannot push the system beyond what the frozen agent can understand and use.
Data efficiency strongly favors tool-centric adaptation. T2 methods like s3 [27] reach competitive or superior performance to A2-style agents such as Search-R1 while using roughly $70\times$ fewer labeled examples, because the T2 subagent only learns a narrow procedural skill (e.g., search policy) rather than relearning general reasoning [27, 108]. Generalization exhibits a similar pattern: T1 tools, by construction, are agent-agnostic and often robust across tasks and agents; T2 tools inherit the inductive biases of strong frozen agents and thus often transfer better across domains than aggressively fine-tuned A1/A2 agents, which may overfit and forget.
Finally, modularity is where T1/T2 shine. Adding a new retrieval strategy, memory writer, or planner simply means training (or swapping in) another tool, leaving the core agent untouched. In A1/A2, the only way to change behavior is to re-adapt the monolithic agent, risking interference with prior skills. In large-scale multi-tool systems, this difference in “evolution ergonomics” is often more decisive than raw performance.

### v3

@@ 5.2.3 Agentic Memory and Skills
The training signal is binary task success or failure: the sparse, trajectory-level reward is broadcast to all case-selection decisions in that trajectory, and a soft Q-learning algorithm updates the retrieval policy. The frozen LLM never sees the Q-values or policy internals; it simply receives retrieved cases as context and generates its plan. Memento achieves top-tier performance: 87.88% on GAIA validation (ranked 1st), 79.40% on GAIA test (3rd place), and 95.0% on SimpleQA. Ablations show that case-based memory adds 4.7–9.6% absolute improvement on out-of-distribution tasks. Only the memory is trained; the same frozen LLM that performed worse without memory now excels because its information diet has been optimized.
Memory operations as learnable skills.
A recent line of work reframes the memory management problem itself: rather than hand-designing fixed operations (append, retrieve, summarize), the operations themselves become learnable and evolvable. MemSkill [203] introduces a controller-executor-designer loop in which a controller selects which memory skill to apply (e.g., extract key facts, consolidate related entries, prune stale information), an LLM executor generates skill-guided memory content, and a designer module analyzes failure cases to evolve the skill set over time. The skill library starts with a small seed set and grows as the designer identifies recurring failure patterns that existing skills cannot address. On LongMemEval [204], LoCoMo, HotpotQA, and ALFWorld, MemSkill outperforms static memory baselines by learning task-appropriate memory granularity, for instance extracting fine-grained entity attributes for QA tasks while retaining coarser episode summaries for planning tasks. The approach demonstrates that memory adaptation need not be limited to what is stored; how the storage operations themselves are performed can also be optimized under frozen-agent supervision.
Test-time memory curation.
@@
Another prominent example of T2 memory adaptation at inference-time is Dynamic Cheatsheet (DC) [205], a lightweight framework that provides a “persistent, evolving memory” for black-box LMs. The system operates “without modifying their underlying parameters” and requires no gradient-based updates. The framework consists of two core modules: a Solution Generator and a Memory Curator. The Memory Curator is the adaptive T2 tool: it operates without access to ground-truth labels, assessing the correctness and efficiency of solutions autonomously after they are produced by the frozen generator. Based on this self-assessment, the curator updates the memory by storing concise, transferable snippets such as “reusable strategies, code snippets, and general problem-solving insights”, rather than full, uncurated transcripts. ReasoningBank [206] extends this test-time curation concept by creating a memory framework that explicitly distills generalizable reasoning strategies from both successful and self-judged failed experiences. Unlike methods that store raw trajectories or only successful routines, ReasoningBank analyzes failures to extract “crucial preventative lessons”. The curated bank of reasoning strategies is then retrieved to guide the agent in future tasks. The framework also introduces memory-aware test-time scaling, which uses the curated memory to guide a scaled exploration, where the diverse experiences from scaling help forge stronger, more generalizable memories. ReasoningBank demonstrates effectiveness on complex benchmarks like WebArena [207] and SWE-Bench [208].
Skill libraries as adaptive tools.
A particularly important form of T2 memory adaptation is the construction of reusable skill libraries, which instantiate skill-based experiential memory [171]. Rather than storing raw trajectories (case-based) or abstract heuristics (strategy-based), these systems distill successful experiences into modular, composable, and executable capabilities that the frozen agent can invoke in future tasks. The skill representations span a continuum from fine-grained code snippets to standardized APIs and MCP protocols [14].
Table 4 organizes representative skill-library systems along two axes: skill granularity (what is stored—code, heuristics, trajectories, specialist agents, or synthesized tools) and acquisition mechanism (how skills are obtained—demonstration, exploration, reflection, or RL). We discuss representative systems from each cell below; the table provides a concise reference for the remainder.
Table 4: Skill-library design space. Systems are organized by skill granularity (what is stored) and acquisition mechanism (how skills are obtained). Domain abbreviations: MC = Minecraft, CU = Computer Use, Web = Web Navigation, Sci = Scientific Domains, Gen = General/Multi-domain.
Granularity Acquisition Representative Systems Domain Key Distinction
Exploration Voyager, OS-Copilot, CRADLE MC, CU Frozen agent generates & verifies code skills
Executable code RL PAE Web Proposer–Agent–Evaluator RL loop
Reflection EXPEL Gen Extracts rules from success & failure
Abstract heuristics Demonstration Synapse Gen Stores state–action exemplars
Full trajectories Exploration JARVIS-1 MC Multimodal (visual+textual) indexing
Exploration SkillWeaver Web Community-shared API-level skills
API / MCP protocols Reflection CASCADE Sci Validates against domain criteria
Programs (compositional) Reflection ASI [209] Gen Structured programs with loops & conditionals
Demonstration LATM Gen Maker–user cost amortization
Synthesized tools Reflection CREATOR Gen Abstract plan $\to$ tool $\to$ verify loop
Specialist agents Exploration AgentStore CU Registry + meta-controller selection
[…]
JARVIS-1 [213] extends the skill-library paradigm to open-world multi-task settings by combining multimodal memory with a pre-trained action planner. The system maintains a multimodal memory that stores successful task-completion trajectories as (observation, plan, action) tuples. Given a new task, JARVIS-1 retrieves the most relevant past trajectories from memory and uses them as in-context examples for the planner. The memory is indexed by both visual similarity (using CLIP embeddings of game frames) and textual similarity (using descriptions of task goals), enabling cross-modal retrieval. On 200+ Minecraft tasks spanning crafting, combat, and exploration, JARVIS-1 achieves substantially higher success rates than Voyager and other baselines, particularly on long-horizon tasks requiring 10+ sequential sub-goals, where memory-guided planning reduces compounding errors.
EXPEL [214] generalizes this principle beyond game environments: a frozen LLM extracts reusable “insights” (abstract rules and heuristics) from both successful and failed trajectories, storing them in a persistent experience pool. On ALFWorld and WebShop, EXPEL improves task success by 18% and 12% respectively over non-adaptive baselines, demonstrating that experiential skill extraction transfers across task instances.
Synapse [215] combines trajectory-level memory with state-action exemplars, enabling a frozen agent to select from a library of demonstrated skills at each decision step.
OS-Copilot [216] extends the skill-library paradigm to general-purpose computer use. The system maintains a self-improving library where each skill is a Python function with a natural-language docstring. When the agent encounters a novel task, it first searches the library for relevant skills; if none exist, it generates a new skill, executes it, and upon success stores it for future reuse. The self-improvement loop operates without human annotation: the agent’s own execution outcomes (success or failure) determine which skills enter the library. On the GAIA benchmark, OS-Copilot with its accumulated skill library outperforms baselines that lack persistent skill storage, confirming that cross-session skill reuse is a key advantage of the T2 approach.
CRADLE [217] applies skill-centric design to general computer control, maintaining a skill curation module that stores, retrieves, and composes reusable action sequences for GUI interaction. The system decomposes complex computer tasks into sub-goals, matches each sub-goal against the skill library, and executes the retrieved skill or generates a new one. Successful skills are refined through repeated use, creating a self-improving repertoire that grows with deployment experience.
AppAgent [218] demonstrates autonomous skill discovery for smartphone use. The agent interacts with mobile apps, discovers UI element functionalities through trial and error, and stores the learned interaction patterns as reusable procedural knowledge in an external document. When encountering a new app or task, the agent retrieves relevant skills from this document to guide its actions. The exploration-based discovery requires no human demonstration and transfers across applications that share similar UI patterns.
Agent S [219] introduces experience-augmented hierarchical planning for computer use, where the agent maintains a library of learned “experience” consisting of successful task-completion strategies. The system uses a Manager-Worker architecture: the Manager decomposes tasks into sub-goals using retrieved experience, and the Worker executes individual actions. The experience retrieval mechanism matches new sub-goals against the library using both semantic similarity and structural task-graph matching, enabling transfer of learned strategies to structurally similar but previously unseen tasks.
AgentStore [220] addresses a different facet of the skill management problem: integrating heterogeneous specialist agents as reusable skills within a unified system. Rather than building a monolithic agent, AgentStore maintains a registry of specialized agents (each trained for a particular domain or tool), and a meta-controller learns to select and compose these specialists for new tasks. The system supports dynamic registration of new specialist agents, enabling the skill library to grow without retraining the meta-controller.
ExACT [221] combines reflective Monte Carlo Tree Search with exploratory learning to teach agents to systematically explore unfamiliar environments. The agent builds a skill tree through exploration, where each node represents a discovered interaction pattern. Successful exploration trajectories are stored as reusable skills, and the MCTS-guided exploration ensures diverse coverage of the action space. The approach bridges the gap between trial-and-error skill discovery (as in AppAgent) and more structured planning-based approaches. Zhao et al. [222] formalize the problem of agentic skill discovery, proposing a framework in which an LLM agent autonomously identifies, abstracts, and catalogs reusable behavioral patterns from its interaction history, providing a theoretical grounding for the empirical skill-library systems described above.
SkillWeaver [223] demonstrates that web agents can autonomously discover, create, and refine reusable API-level skills through exploration. The agent navigates websites, identifies recurring interaction patterns, and encodes them as callable API functions that can be shared across agents and tasks. The collective skill library grows through a community mechanism in which multiple agents contribute discovered skills, enabling rapid coverage of diverse web domains without centralized supervision. PAE [224] (Proposer-Agent-Evaluator) formalizes autonomous skill discovery as a three-component loop: a Proposer generates candidate tasks, an Agent attempts them, and an Evaluator scores the outcomes to produce reward signals for reinforcement learning. The decoupled design allows each component to be optimized independently, and the RL-based training enables the agent to acquire skills that are difficult to specify through demonstrations alone. On web navigation benchmarks, PAE discovers skills that transfer across websites and task types, achieving competitive performance with substantially less human supervision than imitation-learning baselines. SAGE [225] pushes this RL direction further by treating the skill library itself as part of the learning loop: during sequential rollouts across related tasks, newly generated skills accumulate in the library and are immediately available for subsequent tasks, while a skill-integrated reward encourages both useful skill creation and effective reuse. This is a particularly clean instantiation of T2-style self-improvement because the agent’s future competence changes through the evolving skill substrate rather than through monolithic retraining alone. Wang et al. [209] take a complementary programmatic approach: rather than storing skills as natural-language descriptions or raw code, ASI induces structured programs (with conditionals, loops, and subroutine calls) that capture the compositional structure of agentic tasks. The programmatic representation enables systematic generalization to longer task horizons and novel task compositions that flat skill libraries struggle with. CASCADE [226] extends skill creation to scientific domains, where an LLM agent autonomously develops and evolves a library of computational chemistry skills (e.g., molecular property prediction, reaction pathway search) through cumulative self-improvement. Each skill is validated against domain-specific correctness criteria before entering the library, and the system tracks skill dependencies to enable compositional reuse across multi-step scientific workflows.
A related but distinct form of skill construction is tool creation: rather than storing successful action traces, the agent synthesizes new reusable tools (typically as executable functions) that encapsulate solutions to recurring subproblems. LATM [227] (Large Language Models as Tool Makers, ICLR 2024) formalizes a two-phase protocol in which a “tool maker” LLM creates Python functions from task demonstrations, and a lightweight “tool user” LLM applies these functions to new instances. Once created, a tool can be reused across hundreds of instances without regeneration, amortizing the cost of the expensive maker model. On reasoning benchmarks (GSM8K, MATH, TabMWP), LATM matches the performance of a single large model while reducing per-instance cost by up to 79%, because the tool user need only call the pre-built function rather than re-derive the solution strategy. CREATOR [228] (EMNLP 2023 Findings) disentangles abstract reasoning from concrete implementation: given a problem, the agent first formulates an abstract solution plan, then creates a tool (code function) that implements the plan, and finally applies the tool. By separating “what to do” from “how to do it,” CREATOR enables the agent to correct tool implementations through a verification-and-refinement loop without re-deriving the abstract strategy. On Creation Challenge and MATH benchmarks, CREATOR outperforms both chain-of-thought and direct code-generation baselines, and the created tools transfer to new problem variants. Both LATM and CREATOR demonstrate that skill construction can operate at the tool-creation level: the agent does not merely retrieve past experiences but synthesizes new, reusable abstractions that compress recurring reasoning patterns into callable functions.
These systems share a common architecture: the frozen agent generates experience, a curation mechanism (code verification, self-reflection, or success filtering) selects what to retain, and a retrieval interface makes accumulated skills available for future reasoning. The architecture maps directly onto the T2 paradigm, where the agent remains fixed and the skill library evolves under agent-derived supervision. A key distinction among these systems is the granularity of the stored skill: Voyager and OS-Copilot store executable code, EXPEL stores abstract heuristics, JARVIS-1 stores full trajectories, AgentStore stores entire specialist agents, LATM and CREATOR store synthesized tool functions, and SkillWeaver stores API-level callable functions. The choice of granularity trades off composability (fine-grained code skills compose more flexibly) against transfer breadth (coarse heuristics generalize more broadly across task distributions). A second axis of variation is the acquisition mechanism: demonstration-based (Synapse), exploration-based (AppAgent, ExACT, PAE), reflection-based (EXPEL, Reflexion), and RL-based (PAE, Memento). Recent work on multi-agent procedural memory adds a third axis. LEGOMem [229] introduces modular procedural memory for multi-agent systems, where each agent maintains a local memory of learned workflows and a coordination layer enables agents to share, compose, and specialize procedural knowledge across the team. The modular design allows individual agents to refine their skill sets independently while benefiting from collective experience, a pattern that extends the single-agent skill library to collaborative settings. At the meta-level, ADAS [230] (Automated Design of Agentic Systems) automates the design of agent architectures themselves, using an LLM to iteratively propose, evaluate, and refine agentic building blocks (prompts, tool configurations, control flows). ADAS can be viewed as a second-order skill management process: rather than accumulating task-level skills, the system accumulates reusable architectural patterns that improve performance across diverse benchmarks.
Skills in embodied and robotic settings.
The skill-library paradigm extends naturally to embodied agents, where skills correspond to physical action sequences grounded in sensorimotor experience. SayCan [231] introduced the idea of grounding LLM-generated plans in robotic affordances: the LLM proposes candidate skills in natural language, and a value function trained on real-world robot data scores each skill by its probability of successful execution in the current state. The product of the LLM’s semantic score and the affordance score selects the next skill, ensuring that plans are both semantically reasonable and physically feasible. ProgPrompt [232] takes a programmatic approach, prompting the LLM to generate Python-like programs that compose primitive robot actions (grasp, place, navigate) into multi-step task plans. The skill primitives are pre-defined, but the composition logic is generated dynamically, enabling zero-shot transfer to novel household tasks in VirtualHome. Eureka [233] addresses a different bottleneck in embodied skill learning: reward function design. The system uses an LLM to generate candidate reward functions as code, evaluates them through physics simulation, and iteratively refines the reward based on training statistics. On 29 robotic manipulation and locomotion tasks in IsaacGym, Eureka-generated rewards match or exceed human-designed rewards in 83% of cases, enabling dexterous pen-spinning that was previously unsolved. RoboGen [234] combines LLM-based task generation with automated skill learning: the LLM proposes new tasks, generates simulation environments, and designs reward functions, while a reinforcement learning agent acquires the corresponding motor skills. The self-supervised loop enables open-ended skill accumulation without human task specification. These embodied systems illustrate that agent skills are not limited to digital environments; the same T2 principle (frozen high-level planner, adaptive skill repertoire) applies when skills involve physical actions grounded in sensorimotor feedback.
Memory navigation and interactive reading.
[…]
Adapting the embedding space.
@@
Training uses supervised learning on parallel sequences where ground-truth tool calls are replaced with toolken placeholders. The loss is masked so that only the toolken predictions (and subsequent argument tokens) contribute gradients to $W_{\tau}$. The approach is parameter-efficient: adding 234 tools requires training only $234\times 4096\approx 1M$ parameters (for LLaMA’s 4096-dimensional embeddings), compared to the 13B+ parameters of the full model. ToolkenGPT achieves 73% one-hop accuracy on FuncQA (vs. 57% for ReAct), 75% supervised accuracy on 234-relation KAMEL, and 68% success on VirtualHome with 58 action/object tools. New tools can be added by expanding $W_{\tau}$ and continuing training, without full model retraining.
ToolkenGPT shows that adaptation can occur at the interface layer (the embedding space) rather than the parameter layer (the LLM weights), offering a middle ground between fully frozen T1 systems and fully fine-tuned A1/A2 systems.
Beyond the paradigms above, several recent approaches further extend the tool adaptation framework. These methods introduce new training objectives, modalities, and architectural innovations that broaden the scope of tool adaptation.
@@
Common patterns emerge across T2 methods: lightweight training of small modules (millions of parameters) while keeping LLMs frozen (billions of parameters), semantic exploitation of rich representations (hidden states, token spaces, vocabularies), modality bridging between vision/retrieval/tools and frozen text LLMs, and strong generalization to zero-shot or unseen settings.
@@ 6 Comparison of Adaptation Paradigms
We now compare the four adaptation paradigms: (A1) Agent Adaptation with Tool Execution Signal, (A2) Agent Adaptation with Agent Output Signal, (T1) Agent-Agnostic Tool Adaptation, and (T2) Agent-Supervised Tool Adaptation. We first establish a conceptual framework, then analyze the agent-centric (A1/A2) and tool-centric (T1/T2) paradigms in depth, with special focus on the emergent “subagent-as-tool” and “graduation” concepts, and conclude with a quantitative synthesis of critical trade-offs.
@@ 6.1 A Framework for Comparison
  • Cost and Flexibility: We use “cost” to refer to compute and engineering effort required for adaptation, and “flexibility” to mean how easily the system’s behavior can be reconfigured. A1/A2 provide high parametric flexibility (the entire agent policy can change), whereas T1/T2 provide high system-level flexibility (capabilities can be added, swapped, or composed via tools) but remain bounded by the frozen agent’s intrinsic reasoning power.
  • Data Efficiency: Beyond raw compute, the amount of training data required differs substantially across paradigms. Recent evidence suggests that T2 methods can match or surpass A2-style end-to-end agent training with orders of magnitude less data, by only training small subagents around a frozen backbone.
@@
  • Modularity and System Evolution: This axis focuses on engineering implications: how easily a system can be extended or maintained over time. Tool-centric paradigms (T1/T2) support modular evolution and hot-swapping of components; agent-centric paradigms (A1/A2) tend to be monolithic and may suffer from catastrophic forgetting when adapted repeatedly.
In summary, agent adaptation (A1/A2) rewrites the entire policy in a single model, offering high parametric flexibility at the cost of expensive retraining and potential side-effects on unrelated behaviors. Tool adaptation (T1/T2) attaches specialized tools that can be added or replaced without destabilizing the base agent, bounded by what the frozen agent can understand and use. Data efficiency and generalization both favor tool-centric adaptation (see §6.4 for quantitative evidence), while modularity—the ability to swap tools without retraining the core—is often more decisive than raw performance in practice.